OpenStudy (anonymous):

integral (x^2 e^-x^2) dx

OpenStudy (anonymous):

Integrate by parts once. The resulting integral will be a simple substitution.

OpenStudy (anonymous):

That's not true.

OpenStudy (anonymous):

Ah, missed the minus sign...

OpenStudy (anonymous):

is u=x^2 and dv=?

OpenStudy (anonymous):

It's not the minus sign, is the fact that its \[ e^{x^2} \] it's called a Gaussian integral.

OpenStudy (anonymous):

@blinn13 disregard my suggestion

OpenStudy (anonymous):

oops, sorry, \[ e^{-x^2} \]

OpenStudy (anonymous):

okay

OpenStudy (experimentx):

the integral is not in closed form.

OpenStudy (experimentx):

*cannot be expressed

OpenStudy (anonymous):

What are the limits of integration?

OpenStudy (anonymous):

0 to 4....im trying to find a centroid

OpenStudy (experimentx):

hmm ... i would use calculator

OpenStudy (anonymous):

i have already found the number i need to show work though :/

OpenStudy (experimentx):

numerical methods?

OpenStudy (anonymous):

Well then there's a mistake somewhere because it can't be done by hand.

OpenStudy (anonymous):

i see...is (e^-x)^2 equal to e^-2x or e^-x^2 then?

OpenStudy (experimentx):

e^(-2x) of course

OpenStudy (anonymous):

thats where my problem was then...thank you

OpenStudy (anonymous):

then it just becomes u-sub.

OpenStudy (experimentx):

yes yes

OpenStudy (anonymous):

Ah. Yep, there we go.

OpenStudy (experimentx):

or by parts

OpenStudy (anonymous):

by parts would be better

OpenStudy (anonymous):

so u=x^2, du=2x dx, dv=e^-2x dx, v=-e^-2x/2..right?

OpenStudy (anonymous):

So by parts was right, eh? hehe

OpenStudy (experimentx):

how do you use IE by parts? here we use \[ \int u v dx = u \int v - \int (du/dx \int v ) \]

OpenStudy (anonymous):

i have never seen that before

OpenStudy (experimentx):

u = x^2 v = e^(-2x)

OpenStudy (anonymous):

would i be able to do it with normal integration by parts?

OpenStudy (experimentx):

yes .. twice

OpenStudy (anonymous):

ok ill try that...my previous values for u and dv are correct?

OpenStudy (experimentx):

i know that type of form exist ... but i don't use it often.

OpenStudy (experimentx):

\[ \begin{align*} \int x^2 e^{-2x}dx dx &= x^2 \int e^{-2x}dx - \int \left( \frac{dx^2}{dx} \int e^{-2x}dx\right )dx \\ &= x^2 \frac{e^{-2x}}{-2 } - \int 2x \frac{e^{-2x}}{-2 }dx\\ &= - \frac{x^2 e^{-2x}}{2 } + \int x e^{-2x}dx \\ &= - \frac{x^2 e^{-2x}}{2 } + x \int e^{-2x}dx - \int \left( \frac{dx}{dx}\int e^{-2x} dx\right )dx\\ \end{align*} \]

OpenStudy (anonymous):

we have not gotten that far but i think it is coming up soon

OpenStudy (experimentx):

this is integration by parts

OpenStudy (anonymous):

i learned it a different way then : \[uv -\int\limits vdu\]

OpenStudy (experimentx):

yes yes i know that ... integration by parts is just opposite of product rule in differentiation

OpenStudy (anonymous):

i know that

OpenStudy (anonymous):

thank you for your help!

OpenStudy (experimentx):

\[ d(uv) = u dv + v du \\ \int d(vu) = \int u dv + \int v du \\ uv = \int u dv + \int v du \\ \int v du = uv - \int u dv \] put v=g(x), u = \( \int f(x) dx \) you get \[ \int g d\left( \int f \right ) = g \int f dx - \int \left(\frac{dg}{dx} \int f dx\right ) dx \] I think is is easier to use.

OpenStudy (experimentx):

\[ \int f g dx = g \int f dx - \int \left(\frac{dg}{dx} \int f dx\right ) dx \]